10. If a doubly-ionized oxygen atom (O–2) is accelerated from rest by going through a potential difference of 32 V, what will be the change in its kinetic energy?

a. 16 eV b. 32 eV c. 64 eV d. 51 eV e. none of the above
Answer #1

Using energy conservation. AKE YV 2 = 241.6*10°C V=32v AKE = 2x1.6x109x32 J - 64x166x1019 J AK = 64 eV Option e